# Dynamical mean field approach to associative memory model with non-monotonic transfer functions

### v1

@@ Abstract
The Hopfield associative memory model stores random patterns in synaptic couplings according to Hebb’s rule and retrieves them through gradient descent on an energy function. This conventional setting, where neurons are assumed to have monotonic transfer functions, has been central to understanding associative memory. Morita (1993), however, showed that introducing non-monotonic transfer functions can dramatically enhance retrieval performance. While this phenomenon has been qualitatively examined, a full quantitative theory remains elusive due to the difficulty of analysis in the absence of an underlying energy function. In this work, we apply dynamical mean-field theory to the discrete-time synchronous retrieval dynamics of the non-monotonic model, which succeeds in accurately characterizing its macroscopic dynamical properties. We also derive conditions for retrieval states, and clarify their relation to previous studies. Our results provide new insights into the non-equilibrium retrieval dynamics of associative memory models.
@@ I Introduction
In this work, our objective is to address this difficulty by applying the dynamical mean-field theory (DMFT), also known as the generating function method or the path-integral method [9, 10, 11, 12]. A key advantage of DMFT is that, by representing the trajectories of the dynamics directly in terms of generating functions, it enables direct analysis of the system’s dynamical behavior regardless of whether an energy function exists. The Morita model was originally proposed for continuous-time retrieval dynamics. However, because DMFT requires a self-consistent determination of both the noise correlation function and the response function of the system components, its application to investigating long term behavior of continuous-time dynamics becomes computationally prohibitive [13, 14, 15]. For this reason, in the present study we focus on the discrete-time synchronous dynamics corresponding to Morita’s original continuous-time model. According to our analysis, when the time step is set to $0.1$, the storage capacity, defined as the maximal number of retrievable random patterns $p_{\rm c}$ for a system of $N$ neurons, reaches approximately $p_{\rm c}\simeq 0.36N$. In addition, the evaluation of the basin of attraction agrees very well with numerical experiments for systems as large as $N=4096$. Our analysis also revealed that the success or failure of retrieval manifests in qualitatively distinct behaviors of the system’s feedback coefficients, which correspond to Onsager’s reaction term in equilibrium theory [16, 17]. Further, we derived the conditions that must be satisfied by stationary states and discussed their relationship to equilibrium analyses presented in previous studies [18, 19].
@@ II Associative memory model with non-monotonic transfer functions
Let us consider a set of $p$ random binary patterns $\bm{\xi}^{1},\ldots,\bm{\xi}^{p}\in\{+1,-1\}^{N}$ sampled from the uniform distribution over $\{+1,-1\}^{p\times N}$. In the Hopfield model, these patterns are embedded in the synaptic connections among $N$ neurons according to Hebb’s rule as
@@
where $\delta_{ij}$ denotes the Kroenecker delta. Hopfield showed that when an input pattern sufficiently close to one of the stored patterns is presented, the network dynamics that locally decrease the energy function
$H(\bm{\sigma})=-\sum_{i<j}J_{ij}\sigma_{i}\sigma_{j},$ (2)
where $\sigma_{i}\in\{+1,-1\}$ represents the state of the $i$-th neuron, converge to a binary configuration very close to one of the stored patterns. This convergence corresponds to the retrieval of the stored memory.
@@
$\sigma_{i}^{t+1}={\rm sign}\left(\sum_{j\neq i}J_{ij}\sigma_{j}^{t}\right),$ (3)
where ${\rm sign}(x)=x/|x|$ for $x\neq 0$. This means that each neuron updates its state via the monotonic transfer function ${\rm sign}(a)$ with imput $a_{i}^{t}=\sum_{j\neq i}J_{ij}\sigma_{j}^{t}$. A continuous-time counterpart,
$\left\{\begin{array}[]{l}da_{i}/d\tau=-a_{i}+\sum_{j\neq i}J_{ij}\tanh\left(a_{j}\right),\cr\sigma_{i}={\rm sign}(a_{i}),\end{array}\right$. (6)
locally minimizes
$H_{\rm cont}(\bm{a})=-\sum_{i<j}J_{ij}\tanh(a_{i})\tanh(a_{j})+\sum_{i}\int_{0}^{\tanh(a_{i})}\tanh^{-1}(s)ds,$ (7)
which reduces to the minimization of Eq. (2) for $|a_{i}|\gg 1$
These dynamics constitute the classical framework of associative memory, but their retrieval capacity is limited by the gradient-descent nature of the dynamics, which give rise to spurious memories. Morita (1993) addressed this limitation by replacing $\tanh(u)$ in Eq. (6) with non-monotonic transfer functions $f(a)$ [7], such as
$f(a)=\frac{1-\exp[-ca]}{1+\exp[-ca]}\cdot\frac{1+\kappa\exp[c^{\prime}(|a|-h)]}{1+\exp[c^{\prime}(|a|-h)]},$ (8)
shown in Fig. 1(a). Because of its nonmonotonicity, the dynamics is no longer described by an energy function; however, the storate capacity is enhanced significantly from the conventional value $p_{\rm c}\simeq 0.138N$. Although its detailed analysis has never been performed, he estimated $p_{\rm c}\simeq 0.32N$ based on numerical experiments for $N=1000$. He also showed that most spurious states are eliminated, yielding qualitatively new retrieval dynamics beyond the conventional energy-based model. In [7], the parameters are tuned around $(c,c^{\prime},h,\kappa)=(50,15,0.5,-1)$, but the resulting retrieval performance is robust to these choices.
@@
The absence of an energy function makes a theoretical analysis of the nonmonotonic model challenging. Ref. [8] analyzed a piecewise-linear variant [Fig. 1(b)] using a geometrical argument and showed that the retrieval state remains locally stable and free of spurious states up to $p_{\rm c}\simeq 0.398N$ under a specific condition. However, their theory is restricted to particular forms of $f(a)$ and focuses only on local stability, thus providing no characterization of the full retrieval dynamics or the basin of attraction. Consequently, a comprehensive theoretical understanding of the Morita model remains an open problem.
@@ III DMFT
$\left\{\begin{array}[]{l}a_{i}^{t+1}=a_{i}^{t}+\gamma\left[-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right],\cr\sigma_{i}^{t}={\rm sign}(a_{i}^{t}),\end{array}\right$. (11)
@@
$Z(\tilde{\bm{\theta}})=\mathop{\rm tr}_{\bm{\sigma}^{0},\tilde{\bm{a}}}P(\bm{\sigma}^{0})\prod_{t=0}^{T}\prod_{i=1}^{N}\delta\left[a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right)\right]\exp\left(i\sum_{t=1}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right),$ (12)
@@
$\left.Z(\tilde{\bm{\theta}})\right|_{\tilde{\bm{\theta}}=0}=1,\qquad\mathbb{E}_{\bm{\sigma}^{0},\bm{a}^{0}}[a_{i}^{t}]=\left.\frac{\partial Z(\tilde{\bm{\theta}})}{\partial\theta_{i}^{t}}\right|_{\tilde{\bm{\theta}}=0},\qquad\mathbb{E}_{\bm{\sigma}^{0},\bm{a}^{0}}[a_{i}^{t}a_{j}^{s}]=\left.\frac{\partial^{2}Z(\tilde{\bm{\theta}})}{\partial\theta_{i}^{t}\partial\theta_{j}^{s}}\right|_{\tilde{\bm{\theta}}=0},\ldots$ (13)
@@
$\quad=\frac{1}{(2\pi)^{N}}\int_{-i\infty}^{+i\infty}dh_{i}^{t},\exp\left[-ih_{i}^{t}\left(a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right)\right)\right],$ (15)
@@
$\times\left.\mathbb{E}_{\tilde{\bm{\xi}}}\left[\exp\left(\sum_{\mu=2}^{\alpha N}\sum_{t=0}^{T}\gamma v_{\mu}^{t}u_{\mu}^{t}\right)\right]\times\exp\left(\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\},$ (19)
@@
$\simeq\mathop{\rm tr}_{\bm{\sigma}^{0},\tilde{\bm{a}},\tilde{\bm{h}},\tilde{\bm{\phi}}}\left\{P(\bm{\sigma}^{0})\exp\left[-\sum_{t=0}^{T}\sum_{i=1}^{N}ih_{i}^{t}(a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+m^{t}+\phi_{i}^{t}+\sum_{s=0}^{t-1}\Lambda(t,s)f(a_{i}^{s})\right)\right]\right$. (25)
$\quad\left.\times\frac{1}{(2\pi)^{{\color[rgb]{0,0,0}N(T+1)/2}}\det(C)^{N/2}}\exp\left(-\frac{1}{2}\sum_{i=1}^{N}\sum_{t,s}C^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)\times\exp\left(\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\}$ (26)
$=\mathop{\rm tr}_{\bm{\sigma}^{0},\tilde{\bm{a}},\tilde{\bm{\phi}}}\left\{P(\bm{\sigma}^{0})\prod_{t=0}^{T}\prod_{i=1}^{N}\delta\left[a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+m^{t}+\phi_{i}^{t}+\sum_{s=0}^{t-1}\Lambda(t,s)f(a_{i}^{s})\right)\right]\right$. (27)
$\quad\left.\times\frac{1}{(2\pi)^{N(T+1)/2}\det(C)^{N/2}}\exp\left(-\frac{1}{2}\sum_{i=1}^{N}\sum_{t,s}C^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)\times\exp\left(\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\}$ (28)
where $\tilde{\bm{\phi}}=\{\phi_{i}^{t}\}$. For a matrix $A$, let $A_{t}$ denote its submatrix composed of elements whose row and column indices are less than or equal to $t$. We also denote its $(t,s)$ component as $A(t,s)$. Using these notations, the matrix $C=(C(t,s))$ is constructed by recursively adding one row and one column at each time step, as
with initial conditions $C(0,0)=\alpha$ and $G(0,0)=0$, where $I_{t}$ denotes the identity matrix of size $t\times t$. The lower-triangular matrix $\Lambda=(\Lambda(t,s))$ is then evaluated recursively as
The derivation of Eqs. (24)–(31) is provided in Appendix A.
Equation (24) can be interpreted as the generating function for a set of $N$ discrete-time Langevin equations driven by time-correlated noise and freedback terms:
@@
Eqs. (32) and (33) constitute the main result of the present paper.
The random variable $\phi_{i}^{t}$ represents the pure noise portion of the cross-talk noise, which is uncorrelated with $f(a_{i}^{t})$, and $C(t,s)$ represents its time correlation. On the other hand, $G(t,s)$ are referred to as the response coefficients since they coincide with $N^{-1}\sum_{i=1}^{N}\mathbb{E_{\tilde{\bm{\phi}}}}\left[\partial f(a_{i}^{t})/\partial\phi_{i}^{s}\right]$ for $N\to\infty$ (see Eq. (67)). The feedback coefficients $\Lambda(t,s)$ physically encode the self-feedback effect mediated by the cross-talk noises and the last term of Eq. (16); this mechanism corresponds to Onsager’s reaction term in equilibrium theory [16, 17]. These are computed from Eqs. (29)–(31), which involve no direct interactions among neurons. Consequently, numerical analysis of Eq. (32) can be performed for much larger system sizes $N$ than would be feasible for direct simulations of Eq. (11). Since the saddle-point method used in deriving Eq. (24) yields exact results in the limit $N\to\infty$, the remaining numerical errors in Eq. (32) originate solely from the sampling of the random process $\tilde{\bm{\phi}}$ [10]. This feature makes the present framework highly suitable for accurately investigating the macroscopic behavior of large systems.
On the other hand, evaluating $C(t,s)$ and $\Lambda(t,s)$ requires matrix inversion at each time step, which scales computationally as $O(t^{3})$ up to the $t$-th time step even when tricks based on the Shur complement formula are employed [20]. This makes it practically difficult to carry out numerical analyses for $t\gtrsim 10^{3}$ even with modern computational resources. Accordingly, it remains challenging to directly apply the present method to the continuous-time limit, where $t\to\infty$ and $\gamma\to 0$ are taken while keeping $\tau=\gamma t=O(1)$.

### v2

@@ Abstract
The Hopfield associative memory model stores random patterns in synaptic couplings according to Hebb’s rule and retrieves them through gradient descent on an energy function. This conventional setting, where neurons are assumed to have monotonic transfer functions, has been central to understanding associative memory. Morita (1993, Neural Netw. 6 115), however, showed that introducing non-monotonic transfer functions can dramatically enhance retrieval performance. While this phenomenon has been qualitatively examined, a full quantitative theory remains elusive due to the difficulty of analysis in the absence of an underlying energy function. In this work, we apply dynamical mean-field theory to the discrete-time synchronous retrieval dynamics of the non-monotonic model, which succeeds in accurately characterizing its macroscopic dynamical properties. We also derive conditions for retrieval states, and clarify their relation to previous studies. Our results provide new insights into the non-equilibrium retrieval dynamics of associative memory models.
@@ I Introduction
In this work, our objective is to address this difficulty by applying the dynamical mean-field theory (DMFT), also known as the generating function method or the path-integral method [9, 10, 11, 12]. A key advantage of DMFT is that, by representing the trajectories of the dynamics directly in terms of generating functions, it enables direct analysis of the system’s dynamical behavior regardless of whether an energy function exists. The Morita model was originally proposed for continuous-time retrieval dynamics. However, because DMFT requires a self-consistent determination of both the noise correlation function and the response function of the system components, its application to investigating long term behavior of continuous-time dynamics becomes computationally prohibitive [13, 14, 15]. For this reason, in the present study we focus on the discrete-time synchronous dynamics corresponding to Morita’s original continuous-time model. According to our analysis, when the time step size is set to $0.1$, the storage capacity, defined as the maximal number of retrievable random patterns $p_{\rm c}$ for a system of $N$ neurons, reaches approximately $p_{\rm c}\simeq 0.36N$. In addition, the evaluation of the basin of attraction agrees very well with numerical experiments for systems as large as $N=4096$. Our analysis also revealed that the success or failure of retrieval manifests in qualitatively distinct behaviors of the system’s feedback coefficients, which correspond to Onsager’s reaction term in equilibrium theory [16, 17]. Further, we derived the conditions that must be satisfied by stationary states and discussed their relationship to equilibrium analyses presented in previous studies [18, 19].
@@ II Associative memory model with non-monotonic transfer functions
Let us consider a set of $p$ random binary patterns $\bm{\xi}^{1},\ldots,\bm{\xi}^{p}\in\{+1,-1\}^{N}$ sampled from the uniform distribution over $\{+1,-1\}^{N\times p}$. In the Hopfield model, these patterns are embedded in the synaptic connections among $N$ neurons according to Hebb’s rule as
@@
where $\delta_{ij}$ denotes the Kronecker delta. Hopfield showed that when an input pattern sufficiently close to one of the stored patterns is presented, the network dynamics that locally decrease the energy function
$H(\bm{s})=-\sum_{i<j}J_{ij}s_{i}s_{j},$ (2)
where $s_{i}\in\{+1,-1\}$ represents the state of the $i$-th neuron, converge to a binary configuration very close to one of the stored patterns. This convergence corresponds to the retrieval of the stored memory.
@@
$s_{i}^{t+1}=\mathrm{sign}\!\left(\sum_{j\neq i}J_{ij}s_{j}^{t}\right),$ (3)
where $\mathrm{sign}(x)=x/|x|$ for $x\neq 0$. In this scheme, each neuron updates its state using the monotonic transfer function $\mathrm{sign}(a)$, with input $a_{i}^{t}=\sum_{j\neq i}J_{ij}s_{j}^{t}$. More generally, using a monotonically increasing transfer function $\sigma(x)$ that satisfies $\sigma(-x)=-\sigma(x)$ and $\lim_{x\to\pm\infty}\sigma(x)={\rm sign}(x)$, the dynamics can be extended to a continuous-time formulation:
$\left\{\begin{array}[]{l}\displaystyle\frac{da_{i}}{d\tau}=-a_{i}+\sum_{j\neq i}J_{ij}\sigma(a_{j}),\\[6.0pt] s_{i}=\mathrm{sign}(a_{i}).\end{array}\right$. (6)
This dynamics can also be viewed as a local minimization of the following energy function:
$H_{\rm cont}(\bm{a})$ $=-\sum_{i<j}J_{ij}\sigma(a_{i})\sigma(a_{j})+\sum_{i}\int_{0}^{\sigma(a_{i})}\sigma^{-1}(v)\,dv,$ (7)
which is defined for any choice of the monotonically increasing transfer function $\sigma(x)$. Here, $\sigma^{-1}(v)$ denotes the inverse function of $\sigma(x)$.
These dynamics constitute the classical framework of associative memory, but their retrieval capacity is limited by the gradient-descent nature of the dynamics, which give rise to spurious memories. Morita (1993) addressed this limitation by replacing $\sigma(x)$ in Eq. (6) with non-monotonic transfer functions $f(x)$ [7], such as
$f(x)=\frac{1-\exp[-cx]}{1+\exp[-cx]}\cdot\frac{1+\kappa\exp[c^{\prime}(|x|-h)]}{1+\exp[c^{\prime}(|x|-h)]},$ (8)
shown in figure 1(a). Because of its non-monotonicity, the inverse function $f^{-1}(v)$ cannot be defined for Eq. (8), which makes the dynamics no longer described by an energy function as Eq. (7). However, the storage capacity is enhanced significantly from the conventional value $p_{\rm c}\simeq 0.138N$. Although its detailed analysis has never been performed, he estimated $p_{\rm c}\simeq 0.32N$ based on numerical experiments for $N=1000$. He also showed that most spurious states are eliminated, yielding qualitatively new retrieval dynamics beyond the conventional energy-based model. In [7], the parameters are tuned around $(c,c^{\prime},h,\kappa)=(50,15,0.5,-1)$, but the resulting retrieval performance is robust to these choices.
@@
The absence of an energy function makes a theoretical analysis of the nonmonotonic model challenging. Ref. [8] analyzed a piecewise-linear variant (figure 1(b)) using a geometrical argument and showed that the retrieval state remains locally stable and free of spurious states up to $p_{\rm c}\simeq 0.398N$ under a specific condition. Another study [20] defined a cost function that characterizes the fixed point of the retrieval state for a different type of non-monotonic transfer function and, using equilibrium statistical mechanics, demonstrated that the storage capacity increases up to $p_{\rm c}\simeq 0.41N$. However, these studies are restricted to particular forms of $f(x)$ and focus only on local stability, thus providing no characterization of the full retrieval dynamics or the basin of attraction. Consequently, a comprehensive theoretical understanding of the Morita model remains an open problem.
@@ III DMFT
$\left\{\begin{array}[]{l}a_{i}^{t+1}=a_{i}^{t}+\gamma\left[-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right],\cr s_{i}^{t}={\rm sign}(a_{i}^{t}),\end{array}\right$. (11)
@@
$Z(\tilde{\bm{\theta}})=\mathop{\rm tr}_{\bm{s}^{0},\tilde{\bm{a}}}P(\bm{s}^{0})\prod_{t=0}^{T}\prod_{i=1}^{N}\delta\left[a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right)\right]\exp\left(i\sum_{t=1}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right),$ (12)
@@
$\left.Z(\tilde{\bm{\theta}})\right|_{\tilde{\bm{\theta}}=0}=1,\qquad\mathbb{E}_{\bm{s}^{0},\bm{a}^{0}}[a_{i}^{t}]=\left.\frac{\partial Z(\tilde{\bm{\theta}})}{\partial(i\theta_{i}^{t})}\right|_{\tilde{\bm{\theta}}=0},\qquad\mathbb{E}_{\bm{s}^{0},\bm{a}^{0}}[a_{i}^{t}a_{j}^{s}]=\left.\frac{\partial^{2}Z(\tilde{\bm{\theta}})}{\partial(i\theta_{i}^{t})\partial(i\theta_{j}^{s})}\right|_{\tilde{\bm{\theta}}=0},\ldots$ (13)
@@
$\quad=\frac{1}{2\pi i}\int_{-i\infty}^{+i\infty}dh_{i}^{t}\exp\left[-ih_{i}^{t}\left(a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+\sum_{j\neq i}J_{ij}f(a_{j}^{t})\right)\right)\right],$ (15)
@@
$\times\left.\mathbb{E}_{\tilde{\bm{\xi}}}\left[\exp\left(\sum_{\mu=2}^{\alpha N}\sum_{t=0}^{T}\gamma v_{\mu}^{t}u_{\mu}^{t}\right)\right]\times\exp\left(i\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\},$ (19)
@@
$\simeq\mathop{\rm tr}_{\bm{s}^{0},\tilde{\bm{a}},\tilde{\bm{h}},\tilde{\bm{\phi}}}\left\{P(\bm{s}^{0})\exp\left[-\sum_{t=0}^{T}\sum_{i=1}^{N}ih_{i}^{t}\left(a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+m^{t}+\phi_{i}^{t}+\sum_{s=0}^{t-1}\Lambda(t,s)f(a_{i}^{s})\right)\right)\right]\right$. (25)
$\quad\left.\times\frac{1}{(2\pi)^{{\color[rgb]{0,0,0}N(T+1)/2}}\det(C)^{N/2}}\exp\left(-\frac{1}{2}\sum_{i=1}^{N}\sum_{t,s}C^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)\times\exp\left(i\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\}$ (26)
$=\mathop{\rm tr}_{\bm{s}^{0},\tilde{\bm{a}},\tilde{\bm{\phi}}}\left\{P(\bm{s}^{0})\prod_{t=0}^{T}\prod_{i=1}^{N}\delta\left[a_{i}^{t+1}-a_{i}^{t}-\gamma\left(-a_{i}^{t}+m^{t}+\phi_{i}^{t}+\sum_{s=0}^{t-1}\Lambda(t,s)f(a_{i}^{s})\right)\right]\right$. (27)
$\quad\left.\times\frac{1}{(2\pi)^{N(T+1)/2}\det(C)^{N/2}}\exp\left(-\frac{1}{2}\sum_{i=1}^{N}\sum_{t,s}C^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)\times\exp\left(i\sum_{t=0}^{T+1}\bm{\theta}^{t}\cdot\bm{a}^{t}\right)\right\}$ (28)
where $\tilde{\bm{\phi}}=\{\phi_{i}^{t}\}$. This result can be interpreted as the generating function for a set of $N$ discrete-time Langevin equations driven by time-correlated noise $\phi_{i}^{t}$ and freedback terms:
@@
Here, $\phi_{i}^{t}$ is regarded as the pure noise portion of the cross-talk noise. The feedback coefficients $\Lambda(t,s)$ physically encode the self-feedback effect mediated by the cross-talk noises and the last term of Eq. (16); this mechanism corresponds to Onsager’s reaction term in equilibrium theory [16, 17].
The matrices $C=\left(C(t,s)\right)$ and $\Lambda=\left(\Lambda(t,s)\right)$ are evaluated as follows. For a matrix $A$, let $A_{t}$ denote its submatrix composed of elements whose row and column indices are less than or equal to $t$. Using this notation, the matrix $C$ is constructed by recursively adding one row and one column at each time step, as
with initial conditions $C(0,0)=\alpha$ and $G(0,0)=0$, where $I_{t}$ denotes the identity matrix of size $t\times t$. Quantities $G(t,s)$ are referred to as the response functions since they coincide with $N^{-1}\sum_{i=1}^{N}\mathbb{E_{\tilde{\bm{\phi}}}}\left[\partial f(a_{i}^{t})/\partial\phi_{i}^{s}\right]$ for $N\to\infty$ (see Eq. (67)). Due to the nature of the construction method, the matrix $G=(G(t,s))$ becomes a lower-triangular matrix. The feedback coefficients are then evaluated recursively as
The derivation of Eqs. (24)–(33), which constitute the main result of the present paper, is provided in Appendix A.
Eqs. (29)–(33) involve no direct interactions among neurons. Consequently, numerical analysis based on these equations can be performed for much larger system sizes $N$ than would be feasible for direct simulations of Eq. (11). Since the saddle-point method used in deriving Eq. (24) yields exact results in the limit $N\to\infty$, the remaining numerical errors in Eq. (29) originate solely from the sampling of the random process $\tilde{\bm{\phi}}$ [10]. This feature makes the present framework highly suitable for accurately investigating the macroscopic behavior of large systems.
On the other hand, evaluating $C(t,s)$ and $\Lambda(t,s)$ requires matrix inversion at each time step, which scales computationally as $O(t^{3})$ up to the $t$-th time step even when tricks based on the Shur complement formula are employed [21]. This makes it practically difficult to carry out numerical analyses for $t\gtrsim O(10^{3})$ even with modern computational resources. Accordingly, it remains challenging to directly apply the present method to the continuous-time limit, where $t\to\infty$ and $\gamma\to 0$ are taken while keeping $\tau=\gamma t=O(1)$.